# ChatDBG: An AI-Powered Debugging Assistant

### v1

@@
Abstract
This paper presents ChatDBG, the first AI-powered debugging assistant. ChatDBG integrates large language models (LLMs) to significantly enhance the capabilities and user-friendliness of conventional debuggers. ChatDBG lets programmers engage in a collaborative dialogue with the debugger, allowing them to pose complex questions about program state, perform root cause analysis for crashes or assertion failures, and explore open-ended queries like why is x null?. To handle these queries, ChatDBG grants the LLM autonomy to take the wheel and drive debugging by issuing commands to navigate through stacks and inspect program state; it then reports its findings and yields back control to the programmer. Our ChatDBG prototype integrates with standard debuggers including LLDB, GDB, and WinDBG for native code and Pdb for Python. Our evaluation across a diverse set of code, including C/C++ code with known bugs and a suite of Python code including standalone scripts and Jupyter notebooks, demonstrates that ChatDBG can successfully analyze root causes, explain bugs, and generate accurate fixes for a wide range of real-world errors. For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%. ChatDBG has seen rapid uptake; it has already been downloaded nearly 30,000 times.
11footnotetext: Equal contribution.footnotetext: Work done at the University of Massachusetts Amherst.
I Introduction
Debuggers help programmers identify and fix bugs by letting them investigate program state and navigate program execution. Debuggers for mainstream languages, including GDB [1], LLDB [2], and WinDBG (for C, C++, and Rust), jdb (Java), Pdb (Python), and the Chrome or Firefox debuggers (for JavaScript), generally provide the same functionality. In particular, most debuggers support observing program execution via tracing and reporting when a program reaches a given line or function of source code; interrupting execution and returning control to the debugger when the program reaches a given line or function via breakpoints, when a particular condition is true via conditional breakpoints, or when a variable changes via watchpoints (a.k.a. data breakpoints); inspecting local variables, globals, heap objects, and backtraces of the call stack; resuming program execution line-by-line (single-step) or at the granularity of function calls; and in some debuggers, stepping backward through execution via reverse debugging, also known as time-travel or omniscient debugging.
@@ I Introduction
This paper introduces ChatDBG, the first AI-powered debugger assistant. ChatDBG integrates into and significantly extends the functionality of standard debuggers. ChatDBG builds on the insight that large language models (LLMs), such as OpenAI’s GPT-4 [3], enable a debugger to leverage insights and intuition from many thousands of programs as well as the vast real-world knowledge embedded in LLMs.
A debugger integrated with ChatDBG continues to provide its full range of functionality, but also lets programmers engage in debugging dialogs where they can ask high-level questions like why is x null here? or why isn’t this value what I expected?. The question can be as simple as why? if a program has crashed or failed an assertion. ChatDBG then orchestrates a conversation with the LLM. A key novelty of ChatDBG is that it grants autonomy to the LLM to “take the wheel” while answering the programmer’s queries. Specifically, the LLM issues “function calls” [4] to run commands in the underlying debugger to investigate program state, execute code, or obtain source code. The results of those calls are sent back to the LLM for use while constructing its response. After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages.
Our prototype of ChatDBG integrates into four widely used debuggers: GDB, LLDB, WinDBG, and Pdb. Our evaluation presents a range of case studies demonstrating that ChatDBG improves significantly on existing debuggers. On a suite of unpublished Python scripts and Jupyter notebooks written by undergraduate students, one or two queries is sufficient for ChatDBG to properly diagnose and fix defects 87% of the time, typically at a cost of under \$0.20 USD. ChatDBG is also effective at identifying causes and providing fixes for a range of real-world bugs in C/C++ code.
This paper makes the following contributions: it introduces ChatDBG, the first AI-powered debugger assistant; it describes the implementation of our ChatDBG prototype; and it introduces the “take the wheel” approach to integrating large language models with developer tools. The paper also presents an evaluation of ChatDBG that demonstrates its significant advantages over existing debugger functionality.
This paper is organized as follows. Section II illustrates the use of ChatDBG to debug a program. Section III describes key related work. Section IV describes ChatDBG’s implementation. Section V presents our evaluation, and Section VI concludes with several promising directions for future work.
@@
from datascience import *
from ds101 import *
def make_marble_sample():
    table = Table().read_table(’marble-sample.csv’)
    return table.column(’color’)
def proportion_blue(sample):
    return sample
def resampled_stats(observed_marbles, num_trials):
    stats = bootstrap_statistic(observed_marbles,
                                proportion_blue,
                                num_trials)
    assert len(stats) == num_trials
    return stats
observed_marbles = make_marble_sample()
stats = resampled_stats(observed_marbles, 5)
assert np.isclose(np.mean(stats), 0.7)
Figure 1: An example program containing several bugs. It is supposed to create an array of marble colors, compute the proportions of blue marbles in resamples of that array, and assert that their mean is about 0.7, the proportion for the array.
\$ chatdbg bootstrap.py Traceback (most recent call last):   File "/python3.11/site-packages/ipdb/__main__.py", ...    pdb._run(stdlib_pdb._ScriptTarget(mainpyfile))  File "/python3.11/pdb.py", line 1659, in _run    self.run(target.code)  File "/python3.11/bdb.py", line 600, in run    exec(cmd, globals, locals)  File "<string>", line 1, in <module>  File "./bootstrap.py", line 19, in <module>    stats = resampled_stats(observed_marbles, 5)            ^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^  File "./bootstrap.py", line 15, in resampled_stats    assert len(stats) == num_trials           ^^^^^^^^^^^^^^^^^^^^^^^^AssertionErrorUncaught exception. Entering post mortem debuggingRunning ’cont’ or ’step’ will restart the program (ChatDBG) p num_trials 5 (ChatDBG) Why doesn’t stats have 5 elements? (ChatDBG) p len(stats) 150 (ChatDBG) info bootstrap_statistic bootstrap_statistic(observed_sample,                     compute_statistic, num_trials)    Creates num_trials resamples of the initial    sample. Returns an array of the provided    statistic for those samples.    * observed_sample: the initial sample, as an                       array.    * compute_statistic: a function that takes a                         sample as an array and                         returns the statistic for                         that sample.    * num_trials: the number of bootstrap samples                  to create.You MUST assume that ‘bootstrap_statistic‘ isspecified and implemented correctly. (ChatDBG) p proportion_blue <function proportion_blue at 0x1050b00e0> (ChatDBG) info proportion_blue   8     def proportion_blue(sample):   9         return sample From the information we have, the stats variable in resampled_stats is supposed to have 5 elements, but it actually has 150 elements… The Python debugger output shows a reference to a variable or function called proportion_blue, which is used as the compute_statistic parameter for bootstrap_statistic. Upon looking at the proportion_blue function’s definition, it is found that the function simply returns the input sample without computing a statistical value. This means that for each of the num_trials (which is 5), bootstrap_statistic is adding the entire sample to the stats array without processing the sample to obtain a single statistic value…
@@
To fix this issue, the proportion_blue function must be corrected to perform the intended computation of returning the proportion of blue marbles in the sample: def proportion_blue(sample):    return np.mean(sample == 'B')
Figure 2: A debugging session with ChatDBG. For brevity, we elide repetitive or unimportant parts. ChatDBG catches the assertion failure on line 15 and prompts the user to enter commands (bold and highlighted in gray). Commands issued by ChatDBG when it takes the wheel are highlighted in yellow, and its response is highlighted in blue.
II Overview
This section illustrates ChatDBG’s features and ability to assist in debugging the program in Figure 1. That program is a distillation of real errors encountered by students in an introductory data science lab. It creates an array observed_marbles representing the colors of marbles (red or blue) in a sample stored in a file. It then calls bootstrap_statistic to create same-sized resamples of that array. That function computes a statistic for each resample and returns an array of those statistics. In this case, the statistic is proportion_blue, the proportion of blue marbles. Given a sufficiently large number of trials, the mean of the resamples’ statistics should be close to 0.7, the proportion of blue marbles in the original sample [5].
The program fails the assertion in resampled_stats, and Figure 2 illustrates a debugging session. To try to figure out what went wrong, the user issues the Pdb command p num_trials to view the value of that variable. Continuing debugging with existing tools would likely involve issuing additional commands, examining data files, source code, and examining library documentation. With ChatDBG, the user instead starts a dialog with the debugger, asking why doesn’t stats have 5 elements? While constructing the answer (in blue), the LLM takes the wheel and directly issues debugger commands (yellow). These include standard Pdb commands and a ChatDBG-specific info command for accessing the source code and docstrings for any user-written code, and the docstrings for library code (which we assume is correct and not the root cause of any error).
\$ chatdbg bootstrap.py Traceback (most recent call last):   File "./bootstrap.py", line 21, in <module>    assert np.isclose(np.mean(stats), 0.7)           ^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^AssertionError... (ChatDBG) Why is the mean of stats not close to 0.7? (ChatDBG) p np.mean(stats) 0.7466666666666667 (ChatDBG) info resampled_stats ... (ChatDBG) info make_marble_sample .... (ChatDBG) info bootstrap_statistic ...
@@ II Overview
1. Check the CSV file marble-sample.csv to verify the proportion of blue to red marbles and ensure it is close to 70% blue marbles. 2. Review the implementation of the proportion_blue function… 3. Verify that the proportion_blue function … is available in the current scope or imported correctly.
(ChatDBG) p proportion_blue(observed_marbles) 0.7 (ChatDBG) That value is correct. (ChatDBG) p stats array([0.66666667, 0.8       , 0.63333333,        0.76666667, 0.86666667]) At this point, the most straightforward explanation for the error is that the number of trials is not enough to accurately reflect the expected mean due to random sampling effects. Increasing the number of trials would likely result in an average statistic that is closer to the expected value of 0.7, following the Law of Large Numbers.
@@
Figure 3: A debugging session demonstrating how ChatDBG incorporates real-world knowledge. After correcting proportion_blue in Figure 1, the program fails on line 21 because the mean proportion of blue marbles in the resamples is not the expected value. ChatDBG identifies high variance resulting from the small number of trials as the root cause.
@@
III Related Work
System & Date Single Stack Break- Cond. Source Trace Display Eval. Watch- Reverse Explain Propose Open
Introduced Step Nav. points " Level State Code points Bugs Fixes Queries
DDT [6], 1961
EXDAMS [7], 1969
Mesa [8], 1979 ✓*
DBX [9], 1981 1990
GDB [10], 1986 1991 2009
@@ III Related Work
TABLE I: Debugger features and their dates of introduction. Most key features have been around for decades. By integrating into modern debuggers (GDB, LLDB, Pdb and WinDBG), ChatDBG inherits all of their features while significantly extending them with functionality to explain bugs and their root causes, propose fixes, and answer arbitrary natural-language queries over program state. (An asterisk or year in italics means the feature is limited in functionality, performance, or depends on specific hardware support; for example, WinDBG’s reverse debugging only works on specific Intel chips.)
@@
Table I presents an overview of previous interactive debuggers, together with their features and date of introduction. The first interactive debugger, DDT, introduced breakpoints, single-stepping, and stack navigation in 1961 [6]. By 1979, the Mesa debugger had most key features of modern debuggers, including source-level debugging, conditional breakpoints, tracing, and the ability to display runtime state and evaluate code [8]. Arbitrary conditional breakpoints date back at least to 1990 with DBX [9]. Watchpoints were introduced by 1991 and have been in GDB since version 4.0.1 [10]. Reverse debugging it was first implemented in EXDAMS in 1969 [7] is present in widely-used debuggers like GDB (in 2009) [11], and WinDBG (in 2017); the rr debugger, built on top of GDB, also supports reverse debugging on Linux platforms [12]. ChatDBG, by integrating into standard debuggers, inherits and extends their functionality.
@@
Ko and Myers present Whyline, an interactive, trace-based debugger that lets programmers select from a range of queries and identifies (via static and dynamic analysis) a timeline that answers the query [13]. Programmers can only select from those queries presented by Whyline as options. In contrast, ChatDBG permits programmers to pose arbitrary queries that it answers via a dialog with an LLM. Whyline’s use of traces gives it the ability to answer questions that might not be straightforward to answer with the current program state but limits its applicability to relatively short-lived executions.
@@
The goal of program slicing, introduced by Weiser in 1981 [14], is to produce a shorter version of a program limited to the source code that could have led to an error. Program slicing has been extensively studied; Weiser’s paper has been cited over 5,000 times to date. As Section IV-G describes, ChatDBG performs backwards slicing to collect code spread across code cells to facilitate debugging of Jupyter notebooks.
Fault localization seeks to identify the likely location of the root cause of a defect; Wong et al. present a survey with over 400 citations [15]. Unlike previous work, ChatDBG performs fault localization by leveraging LLM-based examination and LLM-driven exploration of source code and program state.
Automated program repair is another highly active area of software engineering research [16]; its goal is to generate source-level program patches that prevent a program from failing. Unlike past work, ChatDBG performs best-effort automated program repair—leveraging its integration within a debugger—via LLM-based examination and LLM-driven exploration of program source and state.
IV Implementation
[…]
1:history = “”
2:loop
3:    line = Input()
4:    if IsDebuggerCommand(line) then
[…]
6:        Print(output)
7:        history = history + (line + $\rightarrow$” + output)
8:    else
[…]
10:             prompt = MakePrompt(Instructions(),
11:                                    EnrichedStack(), Inputs(),
12:                                    Error(), history, line)
13:         else
14:             prompt = MakePrompt(history, line)
[…]
16:         history = “”
17:         while ResponsePending() do
18:             match Receive()
[…]
20:                      output = DoCommand(cmd)
21:                      Print (cmd + $\rightarrow$” + output)
22:                      Send(output)
[…]
Figure 4: ChatDBG architecture and top-level command processing loop.
IV-A Using ChatDBG: Preliminaries

### v2

@@
Abstract.
Debugging is a critical but challenging task for programmers. This paper proposes ChatDBG, an AI-powered debugging assistant. ChatDBG integrates large language models (LLMs) to significantly enhance the capabilities and user-friendliness of conventional debuggers. ChatDBG lets programmers engage in a collaborative dialogue with the debugger, allowing them to pose complex questions about program state, perform root cause analysis for crashes or assertion failures, and explore open-ended queries like ‘why is x null?’. To handle these queries, ChatDBG grants the LLM autonomy to take the wheel: it can act as an independent agent capable of querying and controlling the debugger to navigate through stacks and inspect program state. It then reports its findings and yields back control to the programmer. Our ChatDBG prototype integrates with standard debuggers including LLDB and GDB for native code and Pdb for Python. Our evaluation across a diverse set of code, including C/C++ code with known bugs and a suite of Python code including standalone scripts and Jupyter notebooks, demonstrates that ChatDBG can successfully analyze root causes, explain bugs, and generate accurate fixes for a wide range of real-world errors. For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%. ChatDBG has seen rapid uptake; it has already been downloaded roughly 50,000 times.
Debugging, Artificial Intelligence, Software Engineering
$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debuggingconference: Under Submission; 2025; Someplace, Somewhere
1. Introduction
Debuggers help programmers identify and fix bugs by letting them investigate program state and navigate program execution. Debuggers for mainstream languages, including GDB (Stallman et al., 2011) and LLDB (LLDB, 2010) (for C, C++, and Rust), JDB (for Java), Pdb (for Python), and the Chrome or Firefox debuggers (for JavaScript), generally provide the same functionality. In particular, most debuggers support observing program execution via tracing and reporting when a program reaches a given line or function of source code; interrupting execution and returning control to the debugger when the program reaches a given line or function via breakpoints, when a particular condition is true via conditional breakpoints, or when a variable changes via watchpoints (a.k.a. data breakpoints); inspecting local variables, globals, heap objects, and backtraces of the call stack; and resuming program execution line-by-line (single-step) or at the granularity of function calls.
@@ 1. Introduction
This paper introduces ChatDBG, the first AI-powered debugger assistant. ChatDBG integrates into and significantly extends the functionality of standard debuggers. ChatDBG builds on the insight that large language models (LLMs), such as OpenAI’s GPT-4 (OpenAI, 2023b), enable a debugger to leverage insights and intuition from the vast real-world knowledge embedded in LLMs.
A debugger integrated with ChatDBG continues to provide its full range of functionality but also lets programmers engage in debugging dialogs where they can ask high-level questions like ‘why is x null here?’ or ‘why isn’t this value what I expected?’. The question can be as simple as ‘why?’ if a program has crashed or failed an assertion. To answer such queries, ChatDBG orchestrates a conversation with an LLM. A key novelty of ChatDBG is that it grants autonomy to the LLM to “take the wheel” and act as an independent agent while answering the programmer’s queries. Specifically, the LLM issues “function calls” (OpenAI, 2023a) to run commands in the underlying debugger to investigate program state, execute code, or obtain source code. The results of those calls are sent back to the LLM for use while constructing its response. After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages. Support for the LLM’s agentic control of the debugger is a key technical contribution.
Our prototype of ChatDBG integrates into three widely used debuggers: GDB, LLDB, and Pdb. Our evaluation presents a range of case studies demonstrating that ChatDBG improves significantly on existing debuggers. On a suite of unpublished Python scripts and Jupyter notebooks written by undergraduate students, one or two queries is sufficient for ChatDBG to properly diagnose and fix defects 87% of the time, typically at a cost well under \$0.20 USD. ChatDBG is also effective at identifying causes and providing fixes for a range of real-world bugs in C/C++ code.
This paper makes the following contributions.
  • It introduces ChatDBG, the first AI-powered debugger assistant.
  • It describes the implementation of our ChatDBG prototype.
  • It introduces the “take the wheel” approach to integrating agentic reasoning in large language models with developer tools.
  • It presents an evaluation of ChatDBG that demonstrates its significant advantages over existing debugger functionality.
ChatDBG is broadly applicable to many domains and programming languages, and we expect it to be particularly useful for novice programmers, who often lack the experience to effectively use debuggers. ChatDBG is also useful for experienced programmers, who can augment debugging sessions with ChatDBG’s reasoning capabilities in a natural, conversational, and interactive way.
This paper is organized as follows. Section 2 illustrates the use of ChatDBG to debug a program. Section 3 describes key related work. Section 4 describes ChatDBG’s implementation. Section 5 presents our evaluation. Section 6 outlines directions for future work, and Section 7 concludes.
@@
[…]
[fontsize=,xleftmargin=20pt,linenos]pythonfigs/bootstrap.py
Figure 1. An example program containing several bugs. It is supposed to create an array of marble colors, compute the proportions of blue marbles in resamples of that array, and assert that their mean is about 0.7, the proportion for the array.
\$ chatdbg bootstrap.py Traceback (most recent call last):   File "/python3.11/site-packages/ipdb/__main__.py", ...     pdb._run(stdlib_pdb._ScriptTarget(mainpyfile))   File "/python3.11/pdb.py", line 1659, in _run     self.run(target.code)   File "/python3.11/bdb.py", line 600, in run     exec(cmd, globals, locals)   File "<string>", line 1, in <module>   File "./bootstrap.py", line 19, in <module>     stats = resampled_stats(observed_marbles, 5)             ^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^   File "./bootstrap.py", line 15, in resampled_stats     assert len(stats) == num_trials            ^^^^^^^^^^^^^^^^^^^^^^^^ AssertionError Uncaught exception. Entering post mortem debugging Running ’cont’ or ’step’ will restart the program (ChatDBG) p num_trials 5 (ChatDBG) Why doesn’t stats have 5 elements?
(ChatDBG) p len(stats) 150
(ChatDBG) info bootstrap_statistic bootstrap_statistic(observed_sample, compute_statistic, num_trials)     Creates num_trials resamples of the initial sample. Returns an array of the     provided statistic for those samples.     * observed_sample: the initial sample, as an array.     * compute_statistic: a function that takes a sample as an array and returns       the statistic for that sample.     * num_trials: the number of bootstrap samples to create. You MUST assume that ‘bootstrap_statistic‘ is specified and implemented correctly.
(ChatDBG) p proportion_blue <function proportion_blue at 0x1050b00e0>
(ChatDBG) info proportion_blue   8     def proportion_blue(sample):   9         return sample
{adjustwidth}
.5em1em From the information we have, the stats variable in resampled_stats is supposed to have 5 elements, but it actually has 150 elements…
The Python debugger output shows a reference to a variable or function called proportion_blue, which is used as the compute_statistic parameter for bootstrap_statistic. Upon looking at the proportion_blue function’s definition, it is found that the function simply returns the input sample without computing a statistical value. This means that for each of the num_trials (which is 5), bootstrap_statistic is adding the entire sample to the stats array without processing the sample to obtain a single statistic value…
@@
To fix this issue, the proportion_blue function must be corrected to perform the intended computation of returning the proportion of blue marbles in the sample:
def proportion_blue(sample):    return np.mean(sample == 'B')
Figure 2. A debugging session with ChatDBG. For brevity, we elide repetitive or unimportant parts. ChatDBG catches the assertion failure on line 15 and prompts the user to enter commands (bold and highlighted in gray). Commands issued by ChatDBG when it takes the wheel are highlighted in yellow, and its response is highlighted in blue.
2. Overview
This section illustrates ChatDBG’s features and ability to assist in debugging the program in Figure 1. That program is a distillation of real errors encountered by students in an introductory data science lab. It creates an array observed_marbles representing the colors of marbles (red or blue) in a sample stored in a file. It then calls bootstrap_statistic to create same-sized resamples of that array. That function computes a statistic for each resample and returns an array of those statistics. In this case, the statistic is proportion_blue, the proportion of blue marbles. Given a sufficiently large number of trials, the mean of the resamples’ statistics should be close to 0.7, the proportion of blue marbles in the original sample (Efron, 1979).
The program fails the assertion in resampled_stats, and Figure 2 illustrates a debugging session. To try to figure out what went wrong, the user issues the Pdb command p num_trials to view the value of that variable. Continuing debugging with existing tools would likely involve issuing additional commands, examining data files, source code, and examining library documentation. With ChatDBG, the user instead starts a dialog with the debugger, asking why doesn’t stats have 5 elements? While constructing the answer (in blue), the LLM takes the wheel and directly issues debugger commands (yellow). These include standard Pdb commands and a ChatDBG-specific info command for accessing the source code and docstrings for any user-written code, as well as the docstrings for library code (which we assume is correct and not the root cause of errors).
\$ chatdbg bootstrap.py Traceback (most recent call last):   File "./bootstrap.py", line 21, in <module>     assert np.isclose(np.mean(stats), 0.7)            ^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^ AssertionError ... (ChatDBG) Why is the mean of stats not close to 0.7?
(ChatDBG) p np.mean(stats) 0.7466666666666667
{adjustwidth}
.5em1em …
(ChatDBG) info resampled_stats ...
(ChatDBG) info make_marble_sample ....
(ChatDBG) info bootstrap_statistic ...
{adjustwidth}
.5em1em
@@ 2. Overview
  1. 1.
    Check the CSV file marble-sample.csv to verify the proportion of blue to red marbles and ensure it is close to 70% blue marbles.
  2. 2.
    Review the implementation of the proportion_blue function…
  3. 3.
    Verify that the proportion_blue function … is available in the current scope or imported correctly.
(ChatDBG) p proportion_blue(observed_marbles) 0.7 (ChatDBG) That value is correct.
{adjustwidth}
.5em1em
(ChatDBG) p stats array([0.66666667, 0.8       , 0.63333333,        0.76666667, 0.86666667])
{adjustwidth}
.5em1em
At this point, the most straightforward explanation for the error is that the number of trials is not enough to accurately reflect the expected mean due to random sampling effects. Increasing the number of trials would likely result in an average statistic that is closer to the expected value of 0.7, following the Law of Large Numbers.
@@
Figure 3. A debugging session demonstrating how ChatDBG incorporates real-world knowledge. After correcting proportion_blue in Figure 1, the program fails on line 21 because the mean proportion of blue marbles in the resamples is not the expected value. ChatDBG identifies high variance resulting from the small number of trials as the root cause.
@@
3. Related Work
System and Date Single Step Stack Navigation Breakpoints (BPs) Conditional BPs Source Level Trace Display State Eval. Code Watchpoints Explain Bugs Propose Fixes
Open Queries
DDT (Kotok, 1961), 1961
EXDAMS (Balzer, 1969), 1969
Mesa (Xerox, Systems Development Department, 1979), 1979 ✓*
Dbx (Linton, 1990), 1981 1990
GDB (Stallman et al., 2011), 1986 1991
@@ 3. Related Work
Table 1. Debugger features and their dates of introduction. Most key features have been around for decades. By integrating into modern debuggers (GDB, LLDB, and Pdb), ChatDBG inherits all of their features while significantly extending them with functionality to explain bugs and their root causes, propose fixes, and answer arbitrary natural-language queries over program state. (An asterisk or year in italics means the feature is limited in functionality, performance, or depends on specific hardware support.)
@@
Table 1 presents an overview of previous interactive debuggers, together with their features. The first interactive debugger, DDT, introduced breakpoints, single-stepping, and stack navigation in 1961 (Kotok, 1961). By 1979, the Mesa debugger had most key features of modern debuggers, including source-level debugging, conditional breakpoints, tracing, and the ability to display run-time state and evaluate code (Xerox, Systems Development Department, 1979). Arbitrary conditional breakpoints date back at least to 1990 with Dbx (Linton, 1990). Watchpoints were introduced by 1991 and have been in GDB since version 3.93.
@@
In other work, Ko and Myers present Whyline, an interactive, trace-based debugger that lets programmers select from a range of queries and identifies (via static and dynamic analysis) a timeline that answers the query (Ko and Myers, 2010). Programmers can only select from those queries presented by Whyline as options. In contrast, ChatDBG permits programmers to pose arbitrary queries that it answers via a dialog with an LLM. Whyline’s use of traces gives it the ability to answer questions that might not be straightforward to answer with the current program state but limits its applicability to relatively short-lived executions.
@@
The goal of program slicing, introduced by Weiser in 1981 (Weiser, 1981), is to produce a shorter version of a program limited to the source code that could have led to an error. Program slicing has been extensively studied; Weiser’s paper has been cited over 5,000 times to date. As Section 4.7 describes, ChatDBG performs backwards slicing to collect code spread across code cells to facilitate debugging of Jupyter notebooks.
Fault localization seeks to identify the likely location of a defect’s root cause. Several prior studies have investigated the use of machine learning and LLMs for fault localization. Some of the studied techniques apply machine learning to source code features, coverage data, or other static code features to predict faulty lines of code, but they do not utilize dynamic state and run-time information. DeepFL (Li et al., 2019), Grace (Lou et al., 2021) and DeepRL4FL (Li et al., 2021) are examples of such systems. Similarly, LLMAO (Yang et al., 2024) employs LLMs to provide suspiciousness scores for each line of code in a given program, but only provides access to the source code. AutoFL (Kang et al., 2024) also utilizes an LLM and enables it to statically retrieve source code and coverage information about the program via function calls. However, the system requires a failing test case as input and does not employ run-time state information.
ChatDBG improves upon these systems by providing an LLM with access to run-time program state and the ability to take control of the underlying debugger. Both of these features enhance the LLM’s ability to provide more accurate and informative feedback to the user. We also note that other fault localization techniques can be used in tandem with ChatDBG to improve results, as suggested by Section 5.1’s utilization of backwards slicing to identify code relevant to a bug in Python notebooks and Section 5.2’s utilization of AddressSanitizer (Serebryany et al., 2012) to provide a better starting point for diagnosing and fixing memory errors in native code.
Automated program repair is another active area of software engineering research (Goues et al., 2019). Systems for automatic program repair attempt to generate source-level program patches that prevent a program from failing. ChatDBG performs best-effort automated program repair by requesting that the LLM propose code fixes as part of its response, ultimately letting the programmer drive code changes using these suggestions. Previous research has shown that automated program repair hints can provide significant help in the debugging process and suggests that the benefits of correct advice outweigh the risk of deceptive ones (Eladawy et al., 2024).
4. Implementation
4.1. Using ChatDBG: Preliminaries